OpenStudy (anonymous):

Totally lost with complex numbers. Stayed home sick so I missed the lesson. The book is confusing. So how do I simplify (4-2i)-(-3+i)

jimthompson5910 (jim_thompson5910):

(4-2i)-(-3+i) 4-2i+3-i (4+3)+(-2i-i) 7-3i So (4-2i)-(-3+i) = 7-3i

OpenStudy (anonymous):

distribute the - in from of the second complex number, and then combine real and imaginary parts =4-2i+3-i =7-3i

OpenStudy (amistre64):

treat "i" as you would any other variable when adding like this

OpenStudy (amistre64):

in fact "i" works out fine if you reat it as another variable; just when simplifying, understand that i^2 = -1

OpenStudy (anonymous):

Kay thanks